OpenStudy (anonymous):

find the integral of xe^(x^2) with respect to x

OpenStudy (solomonzelman):

u=x^2

OpenStudy (anonymous):

hmm i was trying that but it wasnt working out

OpenStudy (solomonzelman):

\[\Large \int\limits_{ }^{ }xe^{x^2}~dx\] \(\large\color{black}{u=x^2 }\) \(\large\color{black}{\frac{du}{dx}(x^2)=2x }\) So, the du = (derivative of u) times dx du = 2x times dx

OpenStudy (solomonzelman):

Your du=2x dx has to modified, since you only got x inside the integral, not to x, so lets go ahead an divide both sides by 2, we get, \(\large\color{black}{ \frac{1}{2}du =x ~dx }\)

OpenStudy (solomonzelman):

So, your subs are: \(\large\color{black}{ u=x^2 }\) and \(\large\color{black}{ \frac{1}{2}du =x ~dx }\) .

OpenStudy (anonymous):

and that would get rid of the x in front of the e^(x^2) right

OpenStudy (kainui):

You might want to just take a minute to take the derivative a similar function if you're unsure of its integral since usually it will follow a similar pattern. \[\LARGE (e^{x^2})'=2xe^{x^2}\] So you can see that if you divide both sides by 2 and integrate both sides you will have the answer to your integral: \[\LARGE \int \frac{1}{2}(e^{x^2})' dx=\int xe^{x^2}dx\] \[\LARGE \frac{1}{2}e^{x^2} +C=\int xe^{x^2}dx\]

OpenStudy (solomonzelman):

yes, @YadielG

OpenStudy (anonymous):

ok awesome thanks because the part that had me confused is that i was doing du/2x instead of keeping the x with the dx

OpenStudy (solomonzelman):

kainui, testing integral by where it came from is a good thing, but I think the user is asked to show the work, not to guess what the derivative that it is coming from....

OpenStudy (solomonzelman):

Yes, so what do you get after substitution?

OpenStudy (anonymous):

(1/2) e^(x^2)

OpenStudy (anonymous):

that should be right

OpenStudy (solomonzelman):

you substitute x^2 with u?

OpenStudy (anonymous):

thank you very much sir

OpenStudy (anonymous):

and yes i did

OpenStudy (solomonzelman):